# Edge-weighted Online Stochastic Matching Under Jaillet-Lu LP

### v1

@@ 4 Algorithm
Is it possible? Consider any pair of offline vertices. Without loss of generality, we can assume there is a second-class online vertex type (with an arrival rate of almost 00) adjacent to them. Intuitively, to make sure that we can match them at the desired rate, we wish that these two offline vertices are matched in a very negatively correlated way, so that there is enough probability that the second-class online vertex has at least one unmatched neighbor. In other words, a necessary condition is, $G$ must be the instance that the pair of offline vertices are matched in the most positively correlated way (with respect to the match rate of each first-class/second-class edge, which is decided by the algorithm).
It’s not hard to see that in Algorithm 2, unfortunately, the two offline vertices are matched in the most negatively correlated way during time $[t_{0},t_{1}]$: When one of them is matched, the other must not be matched by a second-class online vertex. However, during other times, they are matched in the most positively correlated way, as we will see later. [^5] During time $[0,t_{0}]$, they are matched independently, which is also the most positively correlated way given the match rate of each first-class/second-class edge, since the match rate of each second-class edge is 00. So, if we restrict $t_{0}=t_{1}$ to remove the middle phase, we will have the desired property.
@@
If we pick the optimal $t_{0}$, Algorithm 3 is slightly worse than Algorithm 2.
@@
Clearly, $g_{u,v}(t)=g(t)$ when $t\leq t_{0}$. So we start with $g_{u,v}(t_{0})=g(t_{0})$ and only consider $t>t_{0}$. When running Algorithm 3, the probability that exactly one offline vertex is unmatched at time $t$ is $2f(t)-g(t)$. In this case, the other one will be matched by both neighbors. So
@@
Notice that the above functions and their derivatives are all continuous in $[t_{0},1]$. Let $t^{\prime}\in[t_{0},1]$ be the latest time such that for any $t\in[0,t^{\prime}]$ we have both $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ and $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$. It suffices to show a contradiction for $t^{\prime}<1$. If $t^{\prime}<1$, let $\varepsilon$ be a sufficiently small constant such that, without loss of generality, $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)<\bar{g}(t)$ and $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ for $t\in(t^{\prime},t^{\prime}+\varepsilon]$. Then, since $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ for $t\leq t^{\prime}+\varepsilon$, we have $f_{v}(t)=f(t)$ for $t\leq t^{\prime}+\varepsilon$ by Corollary 8. Combining with Lemma 7, we get $\bar{f}_{u}(t)\geq\bar{f}_{v}(t)$ for $t\leq t^{\prime}+\varepsilon$. Then we have:
@@
Lemma 5 tells us that $\frac{q_{1}(t)}{1-\ln 2}=\frac{q_{2}(t)}{\ln 2}=\Gamma\approx 0.66217$. Combining it with Lemma 14, we know that every edge $(i,j)$ is matched with probability $\Gamma x_{ij}$. Then the expected objective of Algorithm 4 is $\Gamma\sum_{(i,j)\in E}w_{ij}x_{ij}$. Comparing it with the optimal LP value $\sum_{(i,j)\in E}w_{ij}x_{ij}$ proves the theorem. ∎
@@ Acknowledgments
We thank Mikkel Thorup for helpful comments.
@@ A.1 Missing Details in Theorem 4
The expected objective of Algorithm 2 can be computed as follows. Let $f(t)$ denote the probability that exactly one offline vertex has been matched at time $t$ and $g(t)$ denote the probability that both offline vertices have been matched at time $t$. By our matching strategy in Algorithm 2, during time $[0,t_{0}]$, each offline vertex will only be matched if its first-class neighbor arrives. So for $t\in[0,t_{0}]$, we have
@@
When both offline vertices are unmatched, as long as any online vertex arrives, one of the offline vertices will be matched. Note that the probability that both offline vertices are unmatched at time $t$ is simply $(1-f(t)-g(t))$. So

### v2

@@ 4 Algorithm
Is it possible? Consider any pair of offline vertices. Without loss of generality, we can assume there is a second-class online vertex type (with an arrival rate of almost 0) adjacent to them. Intuitively, to make sure that we can match them at the desired rate, we wish that these two offline vertices are matched in a very negatively correlated way, so that there is enough probability that the second-class online vertex has at least one unmatched neighbor. In other words, a necessary condition is, $G$ must be the instance that the pair of offline vertices are matched in the most positively correlated way (with respect to the match rate of each first-class/second-class edge, which is decided by the algorithm).
It’s not hard to see that in Algorithm 2, unfortunately, the two offline vertices are actually matched in the most negatively correlated way during time $t_{0},t_{1}]$: When one of them is matched, the other must not be matched by a second-class online vertex. However, during other times, they are matched in the most positively correlated way, as we will see later. [^6] During time $0,t_{0}]$, they are matched independently, which is also the most positively correlated way given the match rate of each first-class/second-class edge, since the match rate of each second-class edge is 0. So, if we restrict $t_{0}=t_{1}$ to remove the middle phase, we will have the desired property.
@@
If we pick the optimal $t_{0}$, Algorithm 3 is only slightly worse than Algorithm 2.
@@
Clearly, $g_{u,v}(t)=g(t)$ when $t\leq t_{0}$. So we start with $g_{u,v}(t_{0})=g(t_{0})$ and only consider $t>t_{0}$. When running Algorithm 3, the probability that exactly one offline vertex is unmatched at time $t$ is 2f(t)g(t)2f(t)-g(t). In this case, the other one will be matched by both neighbors. So
@@
Notice that the above functions and their derivatives are all continuous in $t_{0},1]$. Let $t^{\prime}\in[t_{0},1]$ be the latest time such that for any $t\in[0,t^{\prime}]$ we have both $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ and $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$. It suffices to show a contradiction for $t^{\prime}<1$. If $t^{\prime}<1$, let $\varepsilon$ be a sufficiently small constant such that, without loss of generality, $2\bar{f}_{u}(t)-g^{\prime}_{u,v}(t)<\bar{g}(t)$ and $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ for $t\in(t^{\prime},t^{\prime}+\varepsilon]$. Then, since $2\bar{f}_{v}(t)-g^{\prime}_{u,v}(t)\geq\bar{g}(t)$ for $t\leq t^{\prime}+\varepsilon$, we have $f_{v}(t)=f(t)$ for $t\leq t^{\prime}+\varepsilon$ by Corollary 8. Combining with Lemma 7, we get $\bar{f}_{u}(t)\geq\bar{f}_{v}(t)$ for $t\leq t^{\prime}+\varepsilon$. Then we have:
@@
Lemma 5 tells us that $\frac{q_{1}(1)}{1-\ln 2}=\frac{q_{2}(1)}{\ln 2}=\Gamma\approx 0.66217$. Combining it with Lemma 14, we know that every edge $(i,j)$ is matched with probability $\Gamma x_{ij}$. Then the expected objective of Algorithm 4 is $\Gamma\sum_{(i,j)\in E}w_{ij}x_{ij}$. Comparing it with the optimal LP value $\sum_{(i,j)\in E}w_{ij}x_{ij}$ proves the theorem. ∎
@@ Acknowledgments
We thank Mikkel Thorup and the anonymous reviewers for helpful comments.
@@ A.1 Missing Details in Theorem 4
The expected objective of Algorithm 2 can be computed as follows. Let $f(t)$ denote the probability that exactly one offline vertex has been matched at time $t$ and $g(t)$ denote the probability that both offline vertices have been matched at time $t$. By our matching strategy in Algorithm 2, during time $0,t_{0}]$, each offline vertex will only be matched if its first-class neighbor arrives. So for $t\in[0,t_{0}]$, we have
@@
When both offline vertices are unmatched, as long as any online vertex arrives, one of the offline vertices will be matched. Note that the probability that both offline vertices are unmatched at time $t$ is simply (1f(t)g(t))(1-f(t)-g(t)). So